# publications.bib

### v3

@@ 6.1. Relationship to Xiao-Zhu
Assume that the basic element $b\in B(G,\mu)_{\mathrm{un}}$ is unramified (See [XZ, Remark 4.2.11] for a classification). Let us look at the middle degree cohomology $H^{d}(R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}})$. By Theorem 1.14, it has a summand isomorphic to
$H^{d}(R\Gamma_{c}(G,b,\mu)\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})).$
To describe this, let $\mathbf{G}^{\prime}$ be the unique $\mathbb{Q}$-inner form of $\mathbf{G}$ such that $\mathbf{G}(\mathbb{A}^{p\infty})\simeq\mathbf{G}^{\prime}(\mathbb{A}^{p\infty})$, $\mathbf{G}^{\prime}(\mathbb{R})$ is compact modulo center, and $\mathbf{G}_{\mathbb{Q}_{p}}\simeq J_{b}$ (See [Han, Proposition 3.1] for the existence). We write $C(\mathbf{G}^{\prime}(\mathbb{Q})\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/K^{p},\overline{\mathbb{F}}_{\ell})$ for the set of all continuous functions on the profinite set $\mathbf{G}^{\prime}(\mathbb{Q})\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/K^{p}$. It is easy to show that one has an isomorphism
@@
for example by combining [Han, Theorem 3.4] and Corollary 3.6. We let $V_{\mu}\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\hat{G})$ be the usual highest weight module of highest weight $\mu$, which in particular agrees with the highest weight tilting module, since $\mu$ is minuscule. We let $b_{T}$ denote the unique (since $b$ is basic) reduction of $b\in B(G)$ to $B(T)$, and regard it as an element in $B(T)\simeq\mathbb{X}^{*}(\hat{T}^{\Gamma})$ in what follows. It should be the case that, under possible additional constraints on $\mathfrak{m}$ depending on $\mu$ (See for example [Ham2, Conjecture 1.25] and [XZ, Definition 1.4.2]), we have an isomorphism
@@
(23) ${}C(K^{p}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{F}}_{\ell})\otimes^{\mathbb{L}}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}\simeq C(K^{p}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}\otimes V_{\mu}|_{\hat{G}^{\Gamma}}(b_{T})[-d](-\frac{d}{2})$
@@
of $G(\mathbb{Q}_{p})$-representations [^8] One should also be able describe the Weil group action, as in [Ham2, Conjecture 1.25]., where we note that $J_{b}\simeq G$ if $b\in B(G,\mu)_{\mathrm{un}}$ since $b$ is basic, and $J_{b}$ must be quasi-split since $b$ is unramified. In particular, by arguing as in Koshikawa [Ko, Page 6], we know that $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ will have irreducible constituents given by the representations of $J_{b}(\mathbb{Q}_{p})$ with Fargues-Scholze parameter equal $\phi_{\mathfrak{m}}$ as conjugacy classes of parameters. Moreover, using that Assumption 4.4 holds for the groups appearing in Table (1), we know by Propsoition 4.5 that they have to be constituents of $i_{B}^{G}(\chi)$, which will also be irreducible under the generic assumption and the constraints appearing in Table (1) (See the proof of Corollary 4.24). Then [Ham2, Conjecture 1.25] would imply that $R\Gamma_{c}(G,b,\mu)[i_{B}^{G}(\chi)]\simeq i_{B}^{G}(\chi)\otimes V_{\mu}|_{\hat{G}^{\Gamma}}(b_{T})[-d](\frac{-d}{2})$ as $G(\mathbb{Q}_{p})$-modules. Assume $\ell$ is banal (i.e coprime to the pro-order of $K_{p}^{\mathrm{hs}}$) then passing to $K_{p}^{\mathrm{hs}}$-invariants, recalling that it is exact under the banal hypothesis, gives us the isomorphism (23). [^9] For this comparison, it would have been more natural to consider an analogue of Theorem 1.14 with $\overline{\mathbb{Q}}_{\ell}$-coefficients. This is indeed doable assuming that $\phi_{\mathfrak{m}}$ admits a $\overline{\mathbb{Z}}_{\ell}$-lattice as in [Ham2, Theorem 1.17]. This integrality condition is however an artifiact of the theory of solid $\overline{\mathbb{Q}}_{\ell}$-sheaves not being properly understood (e.g excision fails) and should be removable with more technology.
@@
If $B(G,\mu)_{\mathrm{un}}$ consists of only the basic element and the $\mu$-ordinary element and $\phi_{T}$ is strongly $\mu$-regular (Definition 4.13) then [Ham2, Conjecture 1.25] is true. In particular, it follows from [Ham2, Theorem 1.27] that the isomorphism (23) can be made unconditional.
We note that this description of the middle degree cohomology on the generic fiber of the Shimura variety at hyperspecial level parallels Theorem [XZ, Theorem 1.14 (1)], describing the middle degree cohomology on the special fiber of the natural integral model.
@@ 6.2. A General Torsion Vanishing Conjecture
Consider now a general Shimura datum $(\mathbf{G},X)$. Let $\Lambda\in\{\overline{\mathbb{Q}}_{\ell},\overline{\mathbb{F}}_{\ell}\}$. If $\Lambda=\overline{\mathbb{F}}_{\ell}$ assume that $\ell$ is very good with respect to $G:=\mathbf{G}_{\mathbb{Q}_{p}}$, as in [FS, Page 33]. We can then look at the $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-representation.
@@
defined by the cohomology at infinite level. By applying Corollary 4.3, we obtain a $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-equivariant decomposition of this
@@
running over semi-simple $L$-parameters $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\Lambda)$. For such a $\phi$, we let $(\phi_{M},M)$ denote a cuspidal support. I.e $M$ is a Levi of $G$ and $\phi_{M}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}M(\Lambda)$ is a supercuspidal $L$-parameter such that $\phi$ is induced by composing with the natural embedding $\phantom{}{}^{L}M(\Lambda)\rightarrow\phantom{}^{L}G(\Lambda)$. We want to describe the degrees of cohomology that $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}$ sits in for suitably nice $\phi$. The case where $\phi$ factors through $M=T$ is covered by Conjecture 1.2. To go beyond this, we give the following definition.
@@
The terminology of "Langlands-Shahidi type" comes from the fact that the representation $r\circ\phi_{M}$ is precisely the representation which appears in the description of the constant term of the usual Eisenstein series via the Langlands-Shahidi method. The motivation for this definition comes from considering the behavior of geometric Eisenstein series over the Fargues-Fontaine curve for general parabolics, by making analogies with the classical theory over function fields, as developed in [BG, Lau]. In particular, this should be the correct definition that guarantees that the eigensheaves $\mathcal{S}_{\phi}$ on $\mathrm{Bun}_{G}$ with eigenvalue $\phi$ are as simple as possible, and the analysis carried out in [Ham2] generalizes to the non-principal case. This is discussed in more detail in [Ham4, Chapter 3]. In addition, we expect that the consequences derived from the analysis in [Ham2] in the principal case should also generalize. More precisely, we conjecture the following generalization of Proposition 4.15 and Corollary 4.24
@@
Let $B(G)_{M}:=\mathrm{Im}(B(M)_{\mathrm{basic}}\rightarrow B(G))$ be the set of $M$-reducible elements, and let $\phi$ be a semi-simple $L$-parameter of Langlands-Shahidi type with cuspidal support $(M,\phi_{M})$. The category $\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\Lambda)_{\phi}$ of $\phi$-local lisse-étale $\Lambda$-sheaves (as defined in Appendix A) breaks up as direct sum
$\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\Lambda)_{\phi}\simeq\bigoplus_{b\in B(G)_{M}}\mathrm{D}(\mathrm{Bun}_{G}^{b},\Lambda)_{\phi}$
via excision, and the !!! and * pushhforwards agree for any smooth irreducible representation $\rho$ of $J_{b}(\mathbb{Q}_{p})$ lying in $\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b},\Lambda)_{\phi}$ for $b\in B(G)_{M}$.
Given a tilting module $V\in\mathrm{Tilt}_{\Lambda}(\phantom{}^{L}G^{I})$, if $\phi$ is of weakly Langlands-Shahidi type then the map induced by associated the Hecke operator
$T_{V}:\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\Lambda)_{\phi}\rightarrow\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\Lambda)^{BW_{\mathbb{Q}_{p}}^{I}}_{\phi}$
is perverse $t$-exact, where the fact the Hecke operator preserves this subcategory is proven as in Lemma 4.2 (2).
@@
During the preparation of this manuscript, Hansen formulated similar conjectures with rational coefficients [HanBeij]. He refers to Langlands-Shahidi parameters as generous parameters [HanBeij, Definition 2.5] and to weakly Langlands-Shahidi parameters as generic semi-simple parameters [HanBeij, Section 2.3]. One can show that these two definitions are equivalent. Indeed, note that the Galois cohomology $H^{1}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ controls the lifts of a semi-simple parameter $\phi_{M}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}M(\Lambda)$ to a $\phantom{}{}^{L}P(\Lambda)$-valued parameter and that such lifts correspond to finding parameters whose semi-simplification is equal to $\phi$. Moreover, insisting that $H^{1}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ is trivial is equivalent to insisting that $R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M})$ is trivial using local Tate-duality and that the Euler-Poincaré characteristic of this complex is 00. This shows the equivalence of the generous condition with the Langlands-Shahidi type condition, using that the stack of Langlands parameters with rational coefficients is reduced. Lastly, the set of such lifts coming from classes in $H^{0}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ will give rise to non Frobenius semi-simple L-parameters allowing one to see that weakly Langlands-Shahidi is equivalent to generic semi-simple.
@@
For $(\mathbf{G},X)$ of PEL type $A$ or $C$, assuming 1.10 and that $\phi$ of Langlands-Shahidi type, we should also obtain a $W_{E_{\mathfrak{p}}}\times G(\mathbb{Q}_{p})$-equivariant direct sum decomposition
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}\simeq\bigoplus_{b\in B(G,\mu)_{M}}(R\Gamma_{c}(G,b,\mu)_{\phi}\otimes^{\mathbb{L}}V_{b})[2d_{b}],$
where $R\Gamma_{c}(G,b,\mu):=\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}},\Lambda(d_{b}))$ and $R\Gamma_{c}(G,b,\mu)_{\phi}$ is the projection applied to the complex viewed as a $G(\mathbb{Q}_{p})$-representation. This should also generalize once one has appropriate general definitions of $\mathrm{Ig}^{b}$ and $\mathrm{Ig}^{b,*}$ so that one can actually define $V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\Lambda)$. Under possible additional constraints on $\phi$, one should also be able to describe the contribution of $R\Gamma_{c}(G,b,\mu)_{\phi}$ in terms of the decomposition $V_{\mu}|_{Z(\hat{M}^{\Gamma})}=\mathcal{T}_{\mu}|_{Z(\hat{M}^{\Gamma})}$ for $b\in B(G)_{M}$ (along the lines of [Ham2, Conjecture 1.25]), as is explained in the toral case in §6.1. It would be interesting to formulate an optimal conjecture.
@@
We believe that this conjecture should be true under just the weakly Langlands-Shahidi condition. However, we strongly suspect that the splitting of the semi-orthogonal decomposition and in turn the splitting of Mantovan’s filtration discussed in the previous Remark should not hold unless the set $B(G,\mu)_{M}$ is a singleton. In particular, in [HanBeij, Section 2.2] Hansen conjectures the existence of perverse sheaves lying $\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\Lambda)_{\phi}$, for which the semi-orthogonal decomposition does not split. Nonetheless, one still expects perverse $t$-exactness of Hecke operators to hold in these cases [HanBeij, Conjecture 2.32].

### v5

@@ 6. Conjectures and Concluding Remarks
In this section, we give some further complements to the main results established in the previous section. In particular, in §6.1, we explain how, by combining Theorem 1.17, with some further results of [24] one can obtain a more precise description of the contribution of a basic unramified element to the cohomology of the torsion cohomology of the global Shimura variety, that is compatible with an analogous description provided by Xiao and Zhu in [67]. In §6.2, we discuss possible generalizations of our results, introducing the notion of a Langlands-Shahidi type parameter, which should be the most general condition for which the analogue of all of our main results hold (Conjectures 6.4, 6.6).
@@ 6.1. Relationship to Xiao-Zhu
We keep the running notation and assumptions of the previous section. In particular, the pair $(\mathbf{G},X)$ will be a Shimura datum of PEL type AC satisfying Assumption 1.13, and $K=K^{p}K_{p}\subset\mathbf{G}(\mathbb{A}_{f})$ will denote a sufficiently small level. We assume that the local group $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ is an unramified group of the form described in (1) with $p$ and $\ell$ satisfying the corresponding conditions and that $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ is a generic maximal ideal in the spherical Hecke algebra of the form described in Theorem 1.11. We will assume that the basic element $b\in B(G,\mu)_{\mathrm{un}}$ is unramified (See [67, Remark 4.2.11] for a classification). Let us look at the middle degree cohomology $H^{d}(R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}})$. By Theorem 1.17, it has a summand isomorphic to
$H^{d}(R\Gamma_{c}(G,b,\mu)_{\phi_{\mathfrak{m}}}\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})).$
To describe this, let $\mathbf{G}^{\prime}$ be the unique $\mathbb{Q}$-inner form of $\mathbf{G}$ such that $\mathbf{G}(\mathbb{A}^{p\infty})\simeq\mathbf{G}^{\prime}(\mathbb{A}^{p\infty})$, $\mathbf{G}^{\prime}(\mathbb{R})$ is compact modulo center, and $\mathbf{G}_{\mathbb{Q}_{p}}\simeq J_{b}$ (See [29, Proposition 3.1] for the existence). We write $C(\mathbf{G}^{\prime}(\mathbb{Q})\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/K^{p},\overline{\mathbb{F}}_{\ell})$ for the set of all continuous functions on the profinite (since $\mathbf{G}^{\prime}_{\mathbb{R}}$ is compact modulo center) set $\mathbf{G}^{\prime}(\mathbb{Q})\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/K^{p}$. It is easy to show that one has an isomorphism
@@
[…]
@@
(24) ${}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}\otimes_{\mathcal{H}(J_{b})}^{\mathbb{L}}C(K^{p}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{F}}_{\ell})$ $\simeq$
$C(K^{p}K_{p}^{\mathrm{hs}}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}\otimes V_{\mu}|_{\hat{G}^{\Gamma}}(b_{T})[-d](-\frac{d}{2})$
@@
of $H_{K_{p}^{\mathrm{hs}}}$-representations [^12] One should also be able describe the Weil group action, as in [24, Conjecture 1.20]. up to taking the semi-simplifications of both sides as $H_{K_{p}^{\mathrm{hs}}}$-representations, where we note that $J_{b}\simeq G$ if $b\in B(G,\mu)_{\mathrm{un}}$ since $b$ is basic, and $J_{b}$ must be quasi-split since $b$ is unramified. Note that, if we were working with $\overline{\mathbb{Q}}_{\ell}$-coefficients [^13] For this comparison, it would have been more natural to consider an analogue of Theorem 1.17 with $\overline{\mathbb{Q}}_{\ell}$-coefficients. This is indeed doable assuming that $\phi_{\mathfrak{m}}$ admits a $\overline{\mathbb{Z}}_{\ell}$-lattice as in [24, Theorem 1.13]. This integrality condition is however an artifact of the theory of solid $\overline{\mathbb{Q}}_{\ell}$-sheaves not being properly understood (e.g excision fails) and should be removable with more technology., the space of automorphic forms $C(K^{p}\backslash\mathbf{G}^{\prime}(\mathbb{A}_{f})/\mathbf{G}^{\prime}(\mathbb{Q}),\overline{\mathbb{Q}}_{\ell})$ with rational coefficients will be semisimple as a $G(\mathbb{Q}_{p})$-representation, by virtue of the fact that $\mathbf{G}^{\prime}_{\mathbb{R}}$ will be compact modulo center, and in this case the semi-simplification would be unnecessary.
In particular, by arguing as in [37, Page 6], we know that $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ will have cohomology sheaves with irreducible constituents given by the representations of $J_{b}(\mathbb{Q}_{p})$ with Fargues-Scholze parameter equal to $\phi_{\mathfrak{m}}$ as conjugacy classes of parameters. Moreover, using that Assumption 4.5 holds for the groups appearing in Table (1), we know by Proposition 4.6 that they have to be constituents of $i_{B}^{G}(\chi)$, which will also be irreducible under the genericity assumption and the constraints appearing in Table (1) (See the proof of Corollary 4.27). In particular, the only irreducible $J_{b}(\mathbb{Q}_{p})=G(\mathbb{Q}_{p})$-representations that will contribute to the $H_{K_{p}^{\mathrm{hs}}}$-semisimplification of the LHS of (24) will be given by $i_{B}^{G}(\chi)$, and [24, Conjecture 1.20] would imply that $R\Gamma_{c}(G,b,\mu)[i_{B}^{G}(\chi)]\simeq i_{B}^{G}(\chi)\otimes V_{\mu}|_{\hat{G}^{\Gamma}}(b_{T})[-d](\frac{-d}{2})$ as $G(\mathbb{Q}_{p})$-modules. If we assume $\ell$ is banal (i.e coprime to the pro-order of $K_{p}^{\mathrm{hs}}$) then passing to $K_{p}^{\mathrm{hs}}$-invariants, recalling that this is exact under the banal hypothesis, gives us the isomorphism (24) after semi-simplification as $H_{K_{p}^{\mathrm{hs}}}$-modules.
@@
If $B(G,\mu)_{\mathrm{un}}$ consists of only the basic element and the $\mu$-ordinary element and $\phi_{T}$ is strongly $\mu$-regular (Definition 4.16) then [24, Conjecture 1.20] is true. In particular, it follows from [24, Theorem 1.21] that the isomorphism (24) can be made unconditional. More precisely, for $(\mathbf{G},X)$ any PEL type datum of type AC satisfying Assumption 1.13 such that the local group $G$ is of the form described in (1), and $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ a generic maximal ideal of the form described in Theorem 1.11 with associated semisimple toral parameter $\phi_{\mathfrak{m}}^{T}$ such that $\phi_{\mathfrak{m}}^{T}$ is strongly $\mu$-regular in the sense of Definition 4.16, then if $B(G,\mu)_{\mathrm{un}}$ consists of only the basic element and the $\mu$-ordinary element the isomorphism (24) holds after semi-simplifying as a $H_{K_{p}^{\mathrm{hs}}}$-module.
We note that this description of the $K_{p}^{\mathrm{hs}}$-invariants of the middle degree cohomology on the generic fiber of the Shimura variety at hyperspecial level parallels Theorem [67, Theorem 1.14 (1)], describing the middle degree cohomology on the special fiber of the natural integral model.
@@ 6.2. A General Torsion Vanishing Conjecture
Consider now a general Shimura datum $(\mathbf{G},X)$. Let $\Lambda\in\{\overline{\mathbb{Q}}_{\ell},\overline{\mathbb{F}}_{\ell}\}$. If $\Lambda=\overline{\mathbb{F}}_{\ell}$ assume that $\ell$ is very decent as in Definition 4.11. We can then look at the $G(\mathbb{Q}_{p})$-representation.
@@
defined by the cohomology at infinite level. By applying Corollary 4.4, we obtain a $G(\mathbb{Q}_{p})$-equivariant decomposition of this
@@
running over semi-simple $L$-parameters $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\Lambda)$. For such a $\phi$, we let $(\phi_{M},M)$ denote a cuspidal support. I.e $M$ is a Levi of $G$ and $\phi_{M}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}M(\Lambda)$ is a supercuspidal $L$-parameter such that $\phi$ is induced (up to $\hat{G}-$conjugacy) by composing with the natural (up to $\hat{G}-$conjugacy) embedding $\phantom{}{}^{L}M(\Lambda)\rightarrow\phantom{}^{L}G(\Lambda)$. We want to describe the degrees of cohomology that $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}$ sits in for suitably nice $\phi$. The case where $\phi$ factors through $M=T$ is covered by Conjecture 1.2. To go beyond this, we give the following definition.
@@
The terminology of "Langlands-Shahidi type" comes from the fact that the representation $r\circ\phi_{M}$ is precisely the representation which appears in the description of the constant terms of the usual Eisenstein series via the Langlands-Shahidi method ([43], [59], [60]). The motivation for this definition comes from considering the behavior of geometric Eisenstein series over the Fargues-Fontaine curve for general parabolics, by making analogies with the classical theory over function fields, as developed in [5, 44]. In particular, this should be the correct definition that guarantees that the eigensheaves $\mathcal{S}_{\phi}$ on $\mathrm{Bun}_{G}$ with eigenvalue $\phi$ are as simple as possible, and the analysis carried out in [24] generalizes to the non-principal case. This is discussed in more detail in [26, Chapter 3]. In addition, we expect that the consequences derived from the analysis in [24] in the principal case should also generalize. More precisely, we conjecture the following generalization of Proposition 4.18 and Corollary 4.27
@@
Let $B(G)_{M}:=\mathrm{Im}(B(M)_{\mathrm{basic}}\rightarrow B(G))$ be the set of $M$-reducible elements, and let $\phi$ be a semi-simple $L$-parameter of Langlands-Shahidi type with cuspidal support $(M,\phi_{M})$. The category $\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\Lambda)_{\phi}$ of $\phi$-local (as defined in Appendix A) lisse-étale ULA $\Lambda$-sheaves breaks up as direct sum
$\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\Lambda)_{\phi}\simeq\bigoplus_{b\in B(G)_{M}}\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G}^{b},\Lambda)_{\phi}$
via excision [^14] We note that a priori there are no excision triangles for the HN-stratification of $\mathrm{Bun}_{G}$, since there is in general no well-behaved !!-pushforward in the lisse formalsim with non-torsion coefficients. However, one can show that such a pushforward exists for the inclusions of Harder-Narasimhan strata, and satisfies the usual excision triangles (See for example the discussion on [33, Page 11]), and so this is well-defined in the present context., and the !! and * pushhforwards agree for any smooth irreducible representation $\rho$ of $J_{b}(\mathbb{Q}_{p})$ lying in $\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b},\Lambda)_{\phi}$ for $b\in B(G)_{M}$.
Given a tilting module $V\in\mathrm{Tilt}_{\Lambda}(\phantom{}^{L}G^{I})$ (where $\mathrm{Tilt}_{\Lambda}(\phantom{}^{L}G^{I})$ is defined as in [24, Section 9]), if $\phi$ is of weakly Langlands-Shahidi type then the map induced by associated the Hecke operator
$T_{V}:\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\Lambda)_{\phi}\rightarrow\mathrm{D}_{\mathrm{lis}}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\Lambda)^{BW_{\mathbb{Q}_{p}}^{I}}_{\phi}$
on the $\phi$-localized ULA subcategory is perverse $t$-exact, where the fact the Hecke operator preserves this subcategory is proven as in Lemma 4.3 (2).
@@
During the preparation of this manuscript, Hansen formulated similar conjectures with rational coefficients [30]. He refers to Langlands-Shahidi parameters as generous parameters [30, Definition 2.5] and to weakly Langlands-Shahidi parameters as generic semi-simple parameters [30, Section 2.3]. One can show that these two definitions are equivalent. Indeed, note that the Galois cohomology $H^{1}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ controls the lifts of a semi-simple parameter $\phi_{M}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}M(\Lambda)$ to a $\phantom{}{}^{L}P(\Lambda)$-valued parameter and that such lifts correspond to finding parameters whose semi-simplification is equal to $\phi$. Moreover, insisting that $H^{1}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ is trivial is equivalent to insisting that $R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M})$ is trivial using local Tate-duality and that the Euler-Poincaré characteristic of this complex is 0. This shows the equivalence of the generous condition with the Langlands-Shahidi type condition, using that the stack of Langlands parameters with rational coefficients is reduced. Lastly, the set of such lifts coming from classes in $H^{0}(R\Gamma(W_{\mathbb{Q}_{p}},r\circ\phi_{M}))$ will give rise to non Frobenius semi-simple L-parameters allowing one to see that weakly Langlands-Shahidi is equivalent to generic semi-simple.
@@
For $(\mathbf{G},X)$ of PEL type AC, assuming 1.13 and that $\phi$ of Langlands-Shahidi type, we should also obtain a $W_{E_{\mathfrak{p}}}\times G(\mathbb{Q}_{p})$-equivariant direct sum decomposition
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\Lambda)_{\phi}\simeq\bigoplus_{b\in B(G,\mu)_{M}}(R\Gamma_{c}(G,b,\mu)_{\phi}\otimes_{\mathcal{H}(J_{b})}^{\mathbb{L}}V_{b})[2d_{b}],$
where $R\Gamma_{c}(G,b,\mu):=\mathop{\rm colim}\nolimits_{K_{p}\rightarrow\{1\}}R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}},\Lambda(d_{b}))$ and $R\Gamma_{c}(G,b,\mu)_{\phi}$ is the projection applied to the complex viewed as a $G(\mathbb{Q}_{p})$-representation. This should also generalize once one has appropriate general definitions of $\mathrm{Ig}^{b}$ and $\mathrm{Ig}^{b,*}$ so that one can actually define $V_{b}:=R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\Lambda)$. Under possible additional constraints on $\phi$, one should also be able to describe the contribution of $R\Gamma_{c}(G,b,\mu)_{\phi}$ in terms of the decomposition $V_{\mu}|_{Z(\hat{M}^{\Gamma})}=\mathcal{T}_{\mu}|_{Z(\hat{M}^{\Gamma})}$ for $b\in B(G)_{M}$ (along the lines of [24, Conjecture 1.20]), as is explained in the toral case in §6.1. It would be interesting to formulate an optimal conjecture.
@@
We believe that this conjecture should be true under just the weakly Langlands-Shahidi condition. However, we strongly suspect that the splitting of the semi-orthogonal decomposition and in turn the splitting of Mantovan’s filtration discussed in the previous Remark should not hold unless the set $B(G,\mu)_{M}$ is a singleton. In particular, in [30, Section 2.2] Hansen conjectures the existence of perverse sheaves lying $\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\Lambda)_{\phi}$, for which the semi-orthogonal decomposition does not split. Nonetheless, one still expects perverse $t$-exactness of Hecke operators to hold in these cases [30, Conjecture 2.32].